### v2

# Interactive-KBQA: Multi-Turn Interactions for Knowledge Base Question Answering with Large Language Models

Abstract
This study explores the realm of knowledge base question answering (KBQA). KBQA is considered a challenging task, particularly in parsing intricate questions into executable logical forms. Traditional semantic parsing (SP)-based methods require extensive data annotations, which result in significant costs. Recently, the advent of few-shot in-context learning, powered by large language models (LLMs), has showcased promising capabilities. However, fully leveraging LLMs to parse questions into logical forms in low-resource scenarios poses a substantial challenge. To tackle these hurdles, we introduce Interactive-KBQA, a framework designed to generate logical forms through direct interaction with knowledge bases (KBs). Within this framework, we have developed three generic APIs for KB interaction. For each category of complex question, we devised exemplars to guide LLMs through the reasoning processes. Our method achieves competitive results on the WebQuestionsSP, ComplexWebQuestions, KQA Pro, and MetaQA datasets with a minimal number of examples (shots). Importantly, our approach supports manual intervention, allowing for the iterative refinement of LLM outputs. By annotating a dataset with step-wise reasoning processes, we showcase our model’s adaptability and highlight its potential for contributing significant enhancements to the field. [^1] Code and data are available at: https://github.com/JimXiongGM/Interactive-KBQA
Interactive-KBQA: Multi-Turn Interactions for Knowledge Base Question Answering with Large Language Models
Guanming Xiong Peking University Beijing, China gm_xiong@pku.edu.cn                        Junwei Baothanks: Corresponding author. Beijing, China baojunwei001@gmail.com                        Wen Zhao Peking University Beijing, China zhaowen@pku.edu.cn
1 Introduction
Figure 1: Overview of the interactive process.
Knowledge base question answering (KBQA) is an increasingly significant research area that leverages structured knowledge bases (KBs) to provide precise answers to natural language (NL) questions. A KBQA system aims to harness the vast, structured information residing in KBs, such as Freebase (Bollacker et al., 2008), Wikidata (Vrandečić and Krötzsch, 2014), or domain-specific databases like the Movie KB (Zhang et al., 2018). Accurately interpreting and responding to user inquiries with data from knowledge bases has potent applications across numerous sectors, making it a focal point of both academic research and industrial innovation.
Recent advancements in KBQA research can generally be classified into two primary approaches: information retrieval (IR)-based methods and semantic parsing (SP)-based methods. IR-based methods focus on understanding the query, retrieving pertinent subgraphs from KBs related to the question, and extracting the answer from these subgraphs (Dong et al., 2023; Zhang et al., 2022a; Yan et al., 2021). Conversely, SP-based approaches convert NL questions into executable logical forms, leveraging pre-trained generative models to interact with the knowledge base and generate answers (Chen et al., 2021a; Ye et al., 2022; Shu et al., 2022; Yu et al., 2023).
The emergence of large language models (LLMs), such as ChatGPT (Ouyang et al., 2022) and GPT-4 (OpenAI, 2023), has opened new avenues for enhancing KBQA systems. These models have demonstrated promising results in reasoning (Wei et al., 2022) and few-shot learning capabilities (Chen, 2023), setting new benchmarks in the KBQA domain (Gu et al., 2023; Li et al., 2023a; Jiang et al., 2023b; Sun et al., 2024; Li et al., 2023b).
Despite these advancements, KBQA systems face several pressing challenges.
Complex query handling. The primary challenge for IR-based approaches is the difficulty in processing complex queries. For instance, questions characterized solely by entity types or concepts, alongside numerical constraints (e.g., "How many basketball players are taller than 2 meters?"), pose significant challenges. Such queries demand an understanding beyond simple entity recognition, rendering them difficult to address with current methodologies.
Resource scarcity for semantic parsing. SP-based approaches require extensive annotated datasets for training, which are resource-intensive to create. This requirement severely limits the scalability of these methods and contributes to the "black box" nature of the reasoning process, which lacks transparency and interpretability.
Underutilization of large language models (LLMs). Despite the demonstrated capabilities of LLMs in reasoning and few-shot learning, existing KBQA approaches have not fully leveraged these strengths. Most current methods either use LLMs as classifiers to identify predicates (Gu et al., 2023; Sun et al., 2024) or as mimics to generate possible logical forms or questions (Li et al., 2023a, b). There remains a significant opportunity to more effectively leverage LLMs in a few-shot setting to enhance the accuracy and interpretability of KBQA systems.
To address these challenges, we introduce the Interactive-KBQA framework, which combines the reasoning power of LLMs with three tools to interact with KBs. By conceptualizing the LLM as an agent and the KB as the environment, Interactive-KBQA facilitates an iterative, dialogue-based problem-solving process. Figure 1 illustrates the overall process. The process adheres to the thought-action paradigm. Given a complex query, the LLM is required to think and provide an action to interact with the KB through a set of specific tools. These tools return the execution results as observations. Specifically, we designed a set of tools that support heterogeneous databases (i.e., Freebase, Wikidata, and a Movie KB) with a unified interaction logic. We categorized complex questions and provided only two annotated exemplars with complete interactive processes for each type, serving as in-context learning demonstrations to prompt the LLM to complete the task. Furthermore, the method introduced in this study allowed for manual intervention. Consequently, we manually annotated a small dataset with a detailed reasoning process, thereby creating a low-resource dataset. Finally, we fine-tuned open-source LLMs on this dataset. The experiments conducted have shown that this method is effective in low-resource contexts. We have released this high-quality dataset to further contribute to NLP research in the community.
The main contributions of this work are summarized as follows:
  • Propose the Interactive-KBQA, a novel framework that harnesses the reasoning capabilities of LLMs for semantic parsing, enabling multi-turn interactions with KBs.
  • Design a unified SPARQL-based toolset and interaction logic that efficiently address a wide array of complex queries.
  • Conduct extensive experiments to demonstrate that our method achieves remarkable performance with few exemplars.
  • Release a human-annotated KBQA dataset with step-wise reasoning processes, serving as a low-resource dataset.
2 Related Work
Knowledge base question answering (KBQA) methods can be broadly classified into two distinct categories: information retrieval (IR)-based methods and semantic parsing (SP)-based methods. These approaches address the challenge of effectively mapping natural language (NL) queries to the structured formats of knowledge bases (KBs).
IR-based methods aim to extract a question-specific subgraph from the KB and employ ranking algorithms to select the top entities or directly generate answers using text decoders (Lan et al., 2023). To bridge the gap between unstructured texts and structured KBs, Dong et al. (2023) proposed a Structured Knowledge-aware Pre-training method to enhance complex subgraph representation learning. Emphasizing the accuracy of subgraph retrieval, Zhang et al. (2022a) developed a trainable, decoupled subgraph retriever that boosts the performance of subgraph-oriented KBQA models. Further, Jiang et al. (2023c) introduced UniKGQA, a unified framework that integrates retrieval and reasoning across architectures and learning parameters. However, IR-based methods typically rely on identifying entities within a query as an initial step for retrieving the relevant subgraph. This reliance poses a notable limitation, especially when handling complex queries.
SP-based methods parse questions into executable logical forms and perform queries against the KB to retrieve answers. Initial works translated questions into intermediate logical forms before execution. Yih et al. (2015) defined a query graph that resembles subgraphs of the KB and can be directly mapped to a logical form. Hu et al. (2018); Lan and Jiang (2020) extended query graphs to include aggregation operators, and Chen et al. (2021b) enhanced candidate query generation by leveraging predictions of query structure.
In recent years, generative models have increasingly been used to directly generate logical forms. Chen et al. (2021a) designed a retriever to fetch relevant KB items and utilized a Grammar-based Decoder based on LSTM for generating S-expressions (Gu et al., 2021). Meanwhile, Das et al. (2021) employed case-based reasoning for the KBQA task, using Big Bird (Zaheer et al., 2020) to generate logical forms by retrieving relevant cases. Notably, works such as Ye et al. (2022), Shu et al. (2022), Zhang et al. (2023), and Yu et al. (2023) involved initially retrieving basic elements from the KB, such as entities, relations, subgraphs, or texts as supplementary information, before directly generating logical forms using the T5 model (Raffel et al., 2020). However, this approach requires substantial amounts of training data and suffers from a lack of transparency in the reasoning process.
Large language models (LLMs) for KBQA leverage the potent few-shot learning capabilities inherent in LLMs. Recent studies have shown that LLMs can significantly enhance reasoning over KBs. By capitalizing on the powerful few-shot learning capabilities, these methods have demonstrated marked improvements in both the accuracy and efficiency of information retrieval from KBs.
In the context of few-shot scenarios, Gu et al. (2023) proposed that LLMs should prioritize evaluating the plausibility of agent plans over directly generating answers. Concurrently, Li et al. (2023a) advocated for the generation of logical forms as initial drafts, which are subsequently refined into executable queries using the KB. From the agent-environment perspective, Jiang et al. (2023b) developed two specialized interfaces for accessing the KB, while Gu et al. (2024) and Liu et al. (2024) designed seven tools to facilitate this process. Sun et al. (2024) introduced a novel approach that enables LLMs to iteratively employ beam search reasoning on a KB. Furthermore, Zong et al. (2024) proposed assigning three distinct roles to the agent for addressing different KBQA subtasks. Jiang et al. (2024) constructed an instruction dataset based on existing KBQA datasets. However, despite employing LLMs, these methods lack the exploration of more complex questions.
3 Approach
Figure 2: An example of the interactive process.  [^3] For brevity, we omit the prefix: PREFIX ns: <http://rdf.freebase.com/ns/>.
3.1 Problem Formulation
This study investigates a semantic parsing (SP) method for knowledge base question answering (KBQA). A knowledge base (KB) is formally represented as $K\in E\times R\times(E\cup L\cup C)$, where $E$ denotes the set of entities, $R$ signifies the set of relations between entities, $C$ represents the set of classes, and $L$ includes the literal values. Given a question $Q$ and a knowledge base $\mathcal{K}$, our goal is to generate an executable SPARQL expression $S$ that aligns with the question. Thus, the task can be formalized as $p(S|Q,\mathcal{K})$.
3.2 Overview
Recent advancements in large language models (LLMs) have demonstrated remarkable capabilities in few-shot learning and reasoning. However, fully leveraging LLMs to tackle complex KBQA challenges remains an elusive goal. To address this gap, we introduce Interactive-KBQA, an interactive method for KBQA that conceptualizes the LLM as an agent and the KB as an environment. This approach enables semantic parsing and SPARQL generation through dialogic interactions. Specifically, we design a unified interaction logic using three generic tools capable of supporting various complex types of questions (e.g., type constraints, count-based queries) across multiple databases (e.g., Freebase, Wikidata, Movie KB). Figure 3 presents an illustrative example of the interactive process.
3.3 Tools for Knowledge Base
The SP-based KBQA method necessitates the identification of elements and the appropriate graph patterns. Hence, it is crucial to design tools at an atomic level to ensure their universality. In line with this principle, we introduce the following three tools.
SearchNodes(name): This function searches for nodes in the KB using the given surface name, name. Its primary purpose is entity linking (EL). Consequently, it not only returns the formal name of a node but also provides distinguishing features of the entity, such as its description and type. Importantly, this tool avoids dataset-specific EL techniques in favor of a generic retrieval approach.
SearchGraphPatterns(sparql, semantic): This function aims to identify and rank graph predicates within the KB that are essential, guided by the semantic parameter. The function requires the input sparql, which should consist of a SPARQL query beginning with "SELECT ?e WHERE ". Following this, it performs a query on the one-hop subgraph centered around the entity ?e. This includes both incoming and outgoing edges. Subsequently, it ranks the retrieved triples based on the semantic relevance to the semantic parameter and the predicates in triples. Ultimately, the tool returns the top K triples. This tool is designed to precisely identify subgraphs while discarding irrelevant information, thereby optimizing the context window’s use in LLMs. It supports flexible operations and is specifically optimized for Compound Value Type [^4] A Compound Value Type is a Type within Freebase designed to represent data where each entry is composed of multiple fields. (CVT) in Freebase by flattening a CVT node to multiple single-hop relationships. For example, to find movies featuring Tom Hanks in Freebase, the usage would be: SearchGraphPatterns(’SELECT ?e WHERE {?e type.object.name "Tom Hanks"@en .}’, semantic="play in film"), and return: [(?e, film.actor.film -> film.performance.film, "Nothing in Common"), ...].
ExecuteSPARQL(sparql): This tool allows for the direct execution of arbitrary SPARQL queries, ensuring unparalleled flexibility.
Implementation details are provided in Appendix A.2.1, and additional usage instructions are described in Appendix A.2.3.
3.4 Interactive Process
Given a question $Q$, we first construct a prompt text:
$\text{Prompt}=\{\text{Inst},E,Q\}$ (1)
where the instruction text, Inst, consists of tool descriptions, tool usage, and the format of interaction. Inst is database-specific and pre-written. $E$ denotes a set of exemplars, and for each type of question, we manually annotate two complete examples in an interactive format.
In each turn $T$, we let the LLM generate an action based on the Prompt and the history $H$ of the interaction. Specifically, this procedure follows:
$a_{T}=\text{LLM}(\{\text{Prompt},H\})$ (2)
$H=\{c_{0},a_{0},o_{0},...,c_{T-1},a_{T-1},o_{T-1}\}$ (3)
where $c$ denotes the intermediate thought process, an action $a$ belongs to the set {SearchNodes, SearchGraphPatterns, ExecuteSPARQL, Done}, and the observation $o$ is determined by executing an action, which is defined as $o_{T}=\text{Tool}(a_{T})$.
We have devised a general thought-action paradigm for KBQA. Thought: Given $Q$, our initial thought $c_{0}$ is to deconstruct it into sub-queries reminiscent of triple forms; for example, the aforementioned example can be decomposed into (Tom Hanks, act in, target movie). $c_{0}$ is not rigidly defined but adopts a free-form approach, leveraging the semantics of the question to facilitate understanding by both LLMs and humans. Except for the first round, the LLM is required to generate a thought $c$ that clearly articulates its reasoning process based on observations. This approach is intended to render the decision-making process explainable. Action: At each turn $T$, the LLM must generate an action $a_{T}$ that concludes the current round, using Python syntax. The tool then parses, executes, and returns the results, serving as the observation [^5] In case of parsing errors, the tool will return a specific error message.. The LLM decides whether to end the dialogue based on the observation. If $a_{T}=\text{Done}$, we output the final observation $o_{T}$ as the answer.
To minimize inference costs, we train a question type classifier on low-resource data and select $E$ based on the predicted type, as further discussed in Appendix A.2.1.
3.5 Solutions for Complex Questions
Interactive-KBQA involves guiding LLMs through reasoning by annotating an interactive inference process. For different types of complex questions, it is crucial to identify patterns, design interaction modes, and label high-quality examples. This section outlines several representative types of complex questions to elucidate the design rationale behind this paper.
For multi-hop questions, our focus at each step is on specific predicates rather than concrete entities. For instance, as depicted in Figure 3, there is no need to locate specific television programs; expressing the graph pattern within SPARQL is sufficient to enable the tool to handle the ranking part of the task. In the case of CVT structures in Freebase, we explicitly describe the reasoning process when encountered. Moreover, we break down the star-shaped CVT structure into multiple one-hop relations and treat them accordingly. For instance, the semantics of the sentence “Tom Hanks plays the role of ‘David Basner’ in the film ‘Nothing in Common’ ” can be represented by two triples: (“Tom Hanks”, film.actor.film $\rightarrow$ film.performance.film, “Nothing in Common”) and (“Tom Hanks”, film.actor.film $\rightarrow$ film.performance.character, “David Basner”). When querying the qualifier structure in Wikidata, which modifies predicates, we devise a specialized SPARQL query pattern and provide a detailed thought process as well. Additional examples are provided in the Appendix A.8.
3.6 Human-Machine Collaborative Annotation
In realistic scenarios, including examples of all question types within the context is impractical for two primary reasons. First, the associated costs are substantial. Second, as discussed by Su et al. (2023); Zhu et al. (2023), LLMs experience a notable decline in performance when input tokens exceed certain limits. Additionally, Lightman et al. (2023) pointed out that process supervision can enhance a model’s generalization capabilities. Consequently, we propose a human-machine collaborative data annotation method. Enabled by the Interactive-KBQA method introduced in this paper, annotating the reasoning process has become more straightforward.
Specifically, when annotators determine that the action $a_{T}$ is unreasonable, it is manually corrected to $a^{\prime}_{T}$. This adjustment is then incorporated into the message to generate the action $a_{T+1}$. Formally,
$a_{T+1}=\text{LLM}(\{\text{Prompt},\{c_{0},a_{0},o_{0},...,c_{T-1},a^{\prime}_{T},o^{\prime}_{T}\}\})$
We set a breakpoint after each round where the LLM generates thoughts and actions, allowing human evaluators to review and decide on their acceptance. Once accepted, the process continues. If rejected, the annotator revises the generated thoughts and actions before proceeding. For each question type, we randomly sampled 50 data points from the training set and manually annotated them to provide low-resource data.
The core principle of our methodology is to emulate the human data annotation process. Annotators intervene in specific scenarios where the model exhibits hallucinations (such as generating predicates not present in the observation), inconsistencies between Thought and Action, or deviations from the correct answer trajectory (for instance, cases where two consecutive rounds are incorrect).
4 Experiment
We examine Interactive-KBQA across a variety of complex question types and diverse databases (DBs).
4.1 Dataset & Preprocessing
Dataset #Type
#Anno
(Train/Test)
#Raw
(Train/Dev/Test)
WebQSP 2 100 / 300 3,098 / - / 1,639
CWQ 4 200 / 600 27,639 / 3,519 / 3,531
KQA Pro 9 450 / 900 94,376 / 11,797 / 11,797
MetaQA 3 150 / 900 329,282 / 39,138 / 30,903
Table 1: Statistics of the datasets.
WebQuestionsSP (WebQSP) (Yih et al., 2016) and ComplexWebQuestions 1.1 (CWQ) (Talmor and Berant, 2018) are extensively used in KBQA research. These datasets comprise natural language questions paired with their corresponding SPARQL queries based on Freebase. Following Chen et al. (2019), we classify WebQSP questions into 1-hop and 2-hop categories based on the length of the inferential relation chain, i.e., the path from the topic entity to the answer node. CWQ extends WebQSP by incorporating four types of complex questions: Conjunction (Conj), Composition (Compo), Comparative (Compa), and Superlative (Super).
KQA Pro (Cao et al., 2022) is a large-scale dataset designed for complex question answering over a dense set of Wikidata entries. It features nine types of complex questions, including Count (Ct), Query Attribute (QA), Query Attribute Qualifier (QAQ), Query Name (QN), Query Relation (QR), Query Relation Qualifier (QRQ), Select Among (SA), Select Between (SB), and Verify (Vf).
MetaQA (Zhang et al., 2018) is built upon a Movie Knowledge Base (KB) and includes three sets of question-answer pairs: 1-hop, 2-hop, and 3-hop. We converted the Movie KB into RDF triples to facilitate querying via SPARQL.
Due to the prohibitive costs associated with utilizing OpenAI, we employed a uniform sampling method across each type to construct a smaller test dataset. For each DB, we sampled 900 instances, distributing these as follows: 150 instances each for WebQSP and CWQ, 100 instances for KQA Pro, and 300 instances for MetaQA. The statistical details of the test dataset and our annotated (Anno) dataset are presented in Table 1. Our sampled dataset presents a balanced composition, enhancing the assessment of models’ capabilities in handling complex questions.
4.2 Baselines
To comprehensively evaluate our approach, we have selected a range of prior state-of-the-art (SOTA) baseline models.
Fine-tuning (FT) on full data. We selected semantic parsing (SP)-based methods as baselines. DeCAF (Yu et al., 2023) generates both logical forms and direct answers to provide the final outcomes. For KQA Pro, we opted for the BART-SPARQL (Cao et al., 2022). For MetaQA, we choose Edge-aware (Zhang et al., 2022b).
Prompting methods. We chose KB-BINDER (Li et al., 2023a) which leverage large language models (LLMs) in a few-shot setting. Additionally, we compared the efficacy of Chain-of-Thought prompting (CoT prompt) (Wei et al., 2022) alongside the technique of Self-Consistency (SC) (Wang et al., 2023), further enriching our analysis. It’s important to note that StructGPT (Jiang et al., 2023b) and ToG (Sun et al., 2024) operate under the assumption that golden entities have been provided. Consequently, the comparison of these methodologies is presented in Section 4.5.
FT on low-resource data. We fine-tuned open-source LLMs (open-LLMs) to directly generate SPARQL queries through supervised fine-tuning (SFT) manner. We reimplemented a prior SOTA approach for each dataset to serve as a comparison.
4.3 Evaluation Metrics
SP-based methods generate logical forms, which consequently produce answers in an unordered manner. These should be evaluated using the F1 score [^6] We use the average of the F1 scores across all instances.. Additionally, we also report the Random Hits@1 (RHits@1) [^7] Randomly selecting an answer for each question 100 times and calculating the average Hits@1. (Shu et al., 2022), and the Exact Match (EM) metric (Talmor and Berant, 2018) for reference. For KQA Pro, we report the accuracy, defined as the condition where the two sets match exactly.
4.4 Results
Method WebQSP CWQ
1-hop 2-hop Overall RHits@1 Conj Compo Compa Super Overall EM
Prior FT SOTA DeCAF $\dagger$ 74.72 76.32 75.52 80.28 69.19 53.54 18.04 28.00 42.19 50.83
[…]
Prompting w/GPT-4 Turbo IO 28.54 50.05 39.29 45.51 47.54 29.71 33.66 24.67 33.89 45.67
CoT 27.85 51.55 39.70 47.52 44.12 26.30 34.39 30.00 33.70 43.67
CoT+SC 26.66 51.35 39.01 47.08 50.65 28.78 36.98 29.78 36.55 47.50
Ours 69.99 72.41 71.20 72.47 47.44 59.00 47.89 41.96 49.07 59.17
Reimplement DeCAF 24.56 27.55 27.55 39.33 32.19 10.16 11.63 6.00 15.00 19.50
Fine-tuning w/open-LLM SFT-SPARQL (7B) 34.39 33.80 34.09 35.68 12.39 20.44 41.10 38.44 28.10 30.00
Ours (7B) 42.02 45.03 43.57 45.09 31.90 30.70 50.98 46.03 39.90 44.00
Ours (13B) 40.68 53.92 54.86 56.25 30.47 34.51 55.98 49.06 42.50 45.67
Method KQA Pro
Ct QA QAQ QN QR QRQ SA SB Vf Overall
Prior FT SOTA BART-SPARQL $\dagger$ 89 92 87 77 95 81 96 94 86 88.56
Prompting w/GPT-4 Turbo IO 27 23 36 40 25 50 11 69 73 39.33
CoT 22 26 35 34 18 46 21 79 77 39.78
CoT+SC 25 28 33 38 22 51 19 86 75 41.89
Ours 74 83 64 73 73 59 80 61 80 71.89
Reimplement BART-SPARQL 37 44 37 36 67 33 49 78 58 48.78
Fine-tuning w/open-LLM SFT-SPARQL (7B) 52 51 52 47 69 37 60 85 67 57.78
Ours (7B) 53 58 69 48 75 48 75 84 70 64.40
Ours (13B) 63 65 55 49 76 52 68 75 62 62.78
Table 2: Results on WebQSP and CWQ. Results tagged with $\dagger$ denote data from original prediction files, but evaluated on consistent test data. Results with $\sharp$ are reprinted from (Li et al., 2023a).
Method MetaQA
1-hop 2-hop 3-hop Overall RHits@1
Edge-aware $\sharp$ 98.50 93.70 91.00 94.40 96.77
KB-BINDER $\ddagger$ 82.15 91.26 99.66 91.02 86.52
Ours w/GPT-4 96.75 98.47 93.55 96.25 95.97
Ours w/SFT (7B) 93.89 85.99 95.61 91.83 91.41
Table 3: Results on MetaQA. The symbol $\sharp$ denotes results reprinted from Zhang et al. (2022b), while $\ddagger$ indicates we reimplemented the results.
Prompting with GPT-4 Turbo
Table 2 and Table 3 present comprehensive comparisons. As indicated in the table, compared to methods that utilize full data, our approach is at a natural disadvantage on the WebQSP and KQA Pro datasets due to the significant difference in the magnitude of training data (~3K/~33K vs. 4-shot/2-shot, respectively). However, on CWQ and MetaQA datasets, our approach overall outperforms baselines. Furthermore, in the cases of comparative and superlative question types, our approach achieves improvements of 29.85% and 13.96%, respectively. This can be attributed to the fact that within the original dataset, these two types of questions each constitute merely 5% (see Table 11). Unlike previous methods, our approach is not constrained by data distribution, rendering it more robust. For conjunction questions in CWQ, CoT+SC and Decaf outperform our method due to redundant constraints inherent in these queries. Such questions require multiple entity constraints (e.g., "What country in the Mediterranean has Zonguldak Province?") yet often only one is needed to pinpoint the answer ("Zonguldak Province" leads to "Turkey", rendering "country in the Mediterranean" extraneous). Our approach, which focuses on identifying each predicate, complicates the response process.
FT with open-LLMs
The results show that our method outperforms all baselines in low-resource scenarios. Specifically, for evaluations involving two question types from CWQ and two types from KQA Pro, the fine-tuned model exceeded the performance of GPT-4 Turbo.
4.5 Impact of Entity Linking (EL)
Dataset Precision Recall F1 MCR
WebQSP 91.50 74.69 80.00 67.42
w/ELQ 93.67 38.69 41.30
CWQ 87.92 70.67 76.06 76.64
w/ELQ 94.36 41.61 43.81
KQA Pro 80.91 75.82 75.35 80.80
MetaQA 97.33 95.89 95.89 100.00
Table 4: Results of entity linking.
Methods WebQSP CWQ
Overall RHits@1 Overall EM
StructGPT $\dagger$ 44.26 - [^8] The output of StructGPT consists of a string that contains answers, making it unsuitable for evaluation with RHits@1. -
ToG $\ddagger$ 36.40 44.80 31.77 41.94
Ours 71.20 72.47 49.07 59.17
Ours w/golden 78.64 79.25 56.74 66.50
Gain 7.44 6.77 7.67 7.33
Table 5: Results on WebQSP and CWQ with golden entities. $\dagger$ indicates original predictions evaluated on a consistent dataset; $\ddagger$ denotes our reimplementation.
The EL performance of our methods [^9] We find the entity surface name rather than mid. (with GPT-4 Turbo) is delineated in Table 4. The table reports a comparison of the results of our method with those of the widely recognized entity linking tool, ELQ (Li et al., 2020), on the WebQSP and CWQ datasets. Moreover, we introduce the Mention Cover Rate (MCR) to quantify the difficulty of EL. MCR is defined as the rate at which the golden entity names appear within the questions. As shown in our analysis, KQA Pro and MetaQA exhibit higher MCR. It is noted that some entity constrains are redundant in KQA Pro [^10] For instance, the answer of ”What is the connection between Steve Jordan (the one whose position is tight end) and Phoenix (the one that is the twinned administrative body of Chengdu)?” is equivalent to that of ”What is the connection between Steve Jordan and Phoenix?”.. Therefore, despite the F1 score being 75.35%, this is not a bottleneck affecting the performance of KQA Pro. Based on these observations, we conducted experiments on the WebQSP and CWQ datasets under the conditions of given golden entities. The results, as presented in Table 5, reveal that our approach significantly outperforms the previous SOTA across both datasets.
4.6 Ablation Study
We perform various ablation studies to understand the importance of different factors in Interactive-KBQA.
Impact of Exemplars
Dataset Precision Recall F1
CWQ 92.76 92.83 92.79
KQA Pro 92.73 92.44 92.42
Table 6: The performances of question type classifiers.
Setting CWQ KQA Pro
F1 AP Acc AP
Ours(cls+2-shot) 54.69 \$0.50 67.78 \$0.38
w/4-shot $\alpha$ 57.19 \$0.70 46.67 \$0.50
w/4-shot $\beta$ 52.22 \$0.55
zero-shot 51.83 \$0.37 25.25 \$0.30
Table 7: The impact of exemplar number and the average price (AP). The KQA Pro configuration $\alpha$ includes QN, QR, QRQ, and Vf, Whereas $\beta$ comprises Ct, QAQ, SA, and SB.
This section conducts an ablation study to examine the influence of exemplars in two representative settings: CWQ, which includes four types and can barely cover all question types within the prompt text, and KQA Pro, which includes nine types and cannot cover all types due to high inference costs.
The performance of classifiers (cls) trained with our annotated dataset is shown in Table 6. Experiments were conducted in 0-shot and 4-shot scenarios on a randomly selected test subset of 100 entries. For CWQ, one exemplar per question type was used, while for KQA Pro, only four types were sampled. In the zero-shot scenario, only the instruction text was used as a prompt. As Table 7 shows, our method’s performance improves with increased question type coverage in prompts. For CWQ 4-shots, we observed a 2.5 point performance increase, but the cost rose by 37.86%, from \$0.5 to \$0.7. KQA Pro results suggest that accurate demonstrations improve performance and reduce costs. For a detailed discussion about the interaction rounds and costs, see Appendix A.5.
Impact of Backbone Model
Model CWQ KQA Pro
OpenAI
GPT-4 Turbo 49.07 71.89
GPT-3.5 Turbo 13.42 47.92
open-source LLM
Mistral-7B FT 39.90 62.24
w/o FT 4.76 20.41
Llama 2 7B FT  [^11] The maximum context of Llama 2, which is 4,096 tokens, is insufficient for direct inference. 30.42 66.33
Llama 2 13B FT 42.50 62.78
Table 8: Performance of different backbone models
As demonstrated in Table 8, GPT-4 significantly outperforms GPT-3.5121212We use gpt-3.5-turbo-1106. in terms of reasoning capabilities. We have also attempted to apply direct reasoning with Mistral 7B. It was observed that the untrained model significantly struggles with complex, multi-turn interactions. This finding underscores the substantial improvements that fine-tuning provides.
4.7 Error Analysis
Error Type WCWQ KQA Pro
Entity Linking 18 7
Predicate Search 6 0
Reasoning Error 32 48
Format Compliance 17 15
Hallucination 19 21
Other 8 9
Table 9: Distribution of error types.
To systemically assess our method’s deficiencies, we first amalgamate WebQSP+CWQ (WCWQ) and randomly select 100 error instances from each for manual inspection. The aggregated statistical findings are detailed in Table 9.
  • Entity Linking Error refers to the failure to locate nodes using the SearchNodes tool. The primary cause of this error is the LLM’s failure to extract the correct entity mention from the question. For example, in the question "Most Anticipated Tour at the Young Hollywood Awards.", the entity should have been identified as "Young Hollywood Award for Most Anticipated Tour", but instead, the LLM mistakenly searched for "Young Hollywood Awards". Besides, In CWQ and WebQSP, the entities in the returned results do not have descriptions, leading the LLM to consider the results incorrect and subsequently re-initiate the search.
  • Predicate Search Error denotes the failure of the SearchGraphPatterns tool to return the necessary information. This error arises when vector search tools fail to return expected results, such as predicates indicating an organization’s headquarters. For instance, the query "locate in" is unable to match the predicate "organization.organization.headquarters".
  • Reasoning Error means that, given the observations, the LLM fails to generate the appropriate SPARQL query. This failure is largely due to an insufficient semantic understanding of the KG schema. Typically, the LLM often fails to understand the CVT structure (e.g. the direction of the predicate), resulting in incomplete or inaccurate SPARQL queries.
  • Format Compliance Error implies that the LLM does not use the tool in the required format. Examples include incorrectly constructed SPARQL queries with improperly formatted time and numerical values, and the introduction of unrecognized additional parameters to the tool.
  • Hallucination includes generating elements that are inconsistent with the observations.
  • Other Error encompasses errors that cannot be categorized under the above types.
More details are presented in the form of case study in the Appendix A.8.
5 Conclusion
Interactive-KBQA introduces a KBQA approach which utilizes an LLM as an agent for performing semantic parsing through multi-round interactions with a KB. Initially, we developed a unified tool and an interaction methodology tailored to various DB schemas. Subsequently, by categorizing complex questions and annotating a minimal set of exemplars, we employed a few-shot learning strategy that guides the LLM to incrementally generate SPARQL queries. Moreover, we introduced a low-resource dataset that demonstrates superior performance when fine-tuned with open-source LLMs.
Limitations
The prompt learning-based approach heavily relies on the capabilities of LLMs, and in scenarios involving multiple rounds of dialogue, the cost of reasoning becomes significantly high. Additionally, it is impractical to adjust the output of LLM when invoking LLM APIs. Therefore, this paper proposes a collaborative human-machine annotation method to mitigate this issue.
Acknowledgments
This work was supported by the National Key Research and Development Program of China under Grant No. 2023YFC3304404. We would also like to thank to Prof. Zhiyuan Liu and Dr. Yujia Qin from THUNLP group for their invaluable suggestions.
Appendix A Appendix
This appendix provides detailed experimental results and offers further discussion.
A.1 Additional Statistics of Datasets & Databases
Database #Node #Rel
Freebase 22,767,149 20,815
Wikidata 17,754 1,267
Movie KG 38,131 9
Table 10: Statistics of the databases.
Question Type Dist. (Train/Dev/Test)
conjunction 42.00 / 43.59 / 44.60
composition 47.27 / 44.76 / 43.78
comparative 5.54 / 6.22 / 6.03
superlative 5.19 / 5.43 / 5.58
Table 11: Distribution (Dist.) of question types in the original CWQ.
Table 10 presents the statistical information for Freebase, Wikidata, and Movie KG. For Freebase, we utilize a script [^13] https://github.com/lanyunshi/Multi-hopComplexKBQA/blob/master/code/FreebaseTool/FilterEnglishTriplets.py to eliminate non-english triples. In the case of Wikidata, we refer to the subset processed by KQA Pro Cao et al. (2022). Meanwhile, Movie KG is derived by converting the knowledge base provided by MetaQA into RDF triples. Additionally, Table 11 outlines the distribution of question types in the original CWQ.
A.2 System Configurations
A.2.1 Implementation Details
In the interactive process, our study invokes the OpenAI GPT4-Turbo (gpt-4-1106-preview) API to serve as a LLM agent. For each type of questions, we selected and annotated two cases with complete interactive processes as exemplars for in-context learning. During the inference stage, for the WebQSP and MetaQA datasets, all annotated demonstrations are employed as few-shot learning examples, specifically, 4-shot for WebQSP and 6-shot for MetaQA. For CWQ and KQA Pro, we train a bert-base-uncased  [^14] https://huggingface.co/bert-base-uncased model based on low-resource training data as a question type classifier (cls), and then select two exemplars based on the predicted question type, namely cls+2-shot for both CWQ and KQA Pro.
In the development of the SearchNodes tool, Elasticsearch [^15] https://github.com/elastic/elasticsearch is employed to extract all node surface names from the Freebase and MetaQA databases, and vector search techniques are implemented to perform queries on nodes within Wikidata. For the ranking algorithm of the SearchGraphPatterns tool, vector retrieval methods are similarly employed. All processes related to vectorization utilize the OpenAI text-embedding-ada-002 API to generate vectors and employ Chroma [^16] https://github.com/chroma-core/chroma for indexing and searching. Moreover, for the functionality of the ExecuteSPARQL tool, Virtuoso [^17] https://github.com/openlink/virtuoso-opensource serves as the underlying graph query engine.
In the process of fine-tuning open-source LLMs, we utilize, Mistral-7B-v0.1 Jiang et al. (2023a) and Llama-2-13b Touvron et al. (2023). For training optimization, DeepSpeed Rasley et al. (2020) is employed, while inference tasks are accelerated using vLLM Kwon et al. (2023). Llama-2-13b is trained on four NVIDIA A100 80GB GPUs, the other experiments are conducted on two NVIDIA A100 80GB GPUs.
A.2.2 Hyper-Parameter Setting
Parameter Value
temperature 0.7
top_p 1
n 6 # 1 when annotating data
[…]
max_tokens 384
Table 12: Assignments of hyper-parameters for inference.
Parameter Value
batch size (per GPU) 1
model max length 4096
learning rate 1e-5
weight decay 0.001
epochs 10
warm up steps 0
gradient accumulation steps 4
zero stage 3
Table 13: Assignments of hyper-parameters for fine-tuning open-source LLMs.
Table 12 presents the parameter configurations for invoking the OpenAI API and fine-tuning open-source LLMs. We employ DeepSpeed Rasley et al. (2020) to efficiently fine-tune open-source LLMs. The hyperparameter settings utilized for training are detailed in Table 13.
For both the SearchNodes and SearchGraphPatterns tools, the number of returned results is set to 10. In the interaction process, the maximum number of turns is limited to 20, corresponding to 10 rounds.
A.2.3 Instruction Text
Instruction text consists of tool descriptions, tool usages and the format of interaction. The tools designed in this work is capable of processing a variety of complex questions through a unified approach on the different databases. Figures 3, 4, and 5 showcase comprehensive instruction texts for Freebase, Wikidata, and Movie KG, respectively. For further insight into the tool’s applications, examples within these figures are highlighted in green.
Figure 3: Instruction text of Freebase.
Figure 4: Instruction text of Wikidata.
Figure 5: Instruction text of Movie KG for MetaQA.
A.3 Additional Results of Entity Linking
WebQSP
Methods 1-hop 2-hop Overall RHits@1
Ours 69.99 72.41 71.20 72.47
Ours w/golden 77.50 79.79 78.64 79.25
Gain 7.51 7.38 7.44 6.77
Table 14: The impact of given golden entities on WebQSP by question type
CWQ
Methods Conj Compo Compa Super Overall EM
Ours 47.44 59.00 47.89 41.96 49.07 59.17
Ours w/golden entity 53.36 68.19 52.73 52.69 56.74 66.50
Gain 5.92 9.19 4.84 10.72 7.67 7.33
Table 15: The impact of given golden entities on CWQ by question type
In Table 14 and Table 15, we report the impact of given golden entities on model performance by question type for WebQSP and CWQ, respectively.
A.4 Additional Results of Exemplars Selection
CWQ
Setting Conj Compo Compa Super Overall Ave Price
Ours (cls+2-shot) 58.64 66.71 41.35 50.79 54.69 \$0.50
w/4-shot 38.02 51.32 75.66 63.37 57.19 \$0.70
zero-shot 75.86 52.92 32.86 44.00 51.83 \$0.37
Table 16: The impact of exemplar selection on CWQ by question type
KQA Pro
Setting Ct QA QAQ QN QR
Ours (cls+2-shot) 50 90 60 60 80
w/4-shot $\alpha$ 20 70 40 40 50
w/4-shot $\beta$ 50 70 50 70 20
zero-shot 18 64 9 55 0
(continued) QRQ SA SB VF Overall Ave Price
Ours (cls+2-shot) 50 70 80 70 67.78 \$0.38
w/4-shot $\alpha$ 50 50 10 90 46.67 \$0.50
w/4-shot $\beta$ 20 80 40 70 52.22 \$0.55
zero-shot 0 45 9 27 25 \$0.30
Table 17: The impact of exemplar selection on KQA Pro by question type
In Table 16 and Table 17, we report the impact of exemplar selection on model performance by question type for CWQ and KQA Pro, respectively. This examination delineates how the choice of exemplars influences the effectiveness of the model across question types, underscoring the significance of tailored exemplar selection in prompt engineering.
The observation shows that the performance of our proposed approach on KQA Pro drops significantly in zero-shot compared to few-shot scenarios as observed in Table 17. This discrepancy arises primarily due to the intricate nature of the questions within the KQA Pro dataset, which involve multi-hop reasoning, constraints, and qualifiers. Constructed using SPARQL templates and phrased naturally by human annotators, these questions are complex. In zero-shot settings, without exemplars, even human annotators may struggle with semantic parsing due to the varied and unseen KG schema. Conversely, the few-shot approach leverages a small set of carefully annotated examples, effectively guiding the parsing process and accounting for the observed performance differences.
A.5 Interaction Turns and Costs
Method WebQSP CWQ
Overall Ave Price Overall Ave Price
Human Anno 0.97 / 10.26 \$0.16 1.34 / 11.59
Ours w/GPT-4 94.33 / 11.01 \$0.39 80.50 / 14.57 \$0.49
Ours w/GPT-4 + Golden 95.33 / 9.17 \$0.24 82.50 / 12.74 \$0.38
Ours w/FT-Mistral 74.05 / 11.71 - 64.67 / 13.29 -
Method KQA Pro MetaQA
Overall Ave Price Overall Ave Price
Human Anno 0.80 / 11.75 \$0.23 0.08 / 7.13 \$0.06
Ours w/GPT-4 93.89 / 11.91 \$0.33 99.67 / 7.45 \$0.14
Ours w/FT-Mistral 81.20 / 11.18 - 98.78 / 7.07 -
Table 18: Comparison of turn count and cost across human-annotated data (Human Anno) and four datasets. For Human Anno, the average number of human interventions and interaction turn are reported. Concerning model performance, the average success rate and interaction turn are reported.
In the process of interaction, the turn constitutes the fundamental unit, with each utterance made by any participant being recorded as a distinct turn. Table 18 provides a statistical analysis of the interaction turns. A case is deemed successful if the LLM agent explicitly generates Action:Done. Otherwise, the dialogue is concluded once the maximum turn number is reached. It is feasible to consider the turn count of human-annotated data as a lower bound. It is observed that, except for on CWQ, GPT-4 have approached this lower bound. Specifically, in CWQ, the overall success rate per turn for GPT-4 is 80.5/14.57, which, compared to WebQSP, shows a decrease of 14.66% and an increase of 39.96% in terms of success rate and turn count, respectively. This also delineates the difficulty level of the dataset from one perspective. In contrast, the KQA Pro dataset exhibits a stronger regularity, hence presenting a relatively lower level of difficulty. Furthermore, we discovered that on complicated question datasets, the average number of turns for fine-tuned LLMs is less than that for GPT-4, indicating that our annotated data effectively reduced the model’s trial-and-error process by providing valuable information. Additionally, we report the average cost per dialogue for reference  [^18] The price of gpt-4-1106-preview: Input \$0.01 / 1K tokens and Output \$0.03 / 1K tokens. Given that we set the return sequence parameter to 6 for inference and to 1 for human annotation, the cost comparison is not strictly equivalent. Complete results categorized by question type are available in the Table 20, Table 21, and Table 22.
Method WebQSP
1-hop 2-hop Overall Ave Price
Human Anno 1.08 / 10.04 0.86 / 10.48 0.97 / 10.26 \$0.16
Ours w/GPT-4 94.67 / 10.83 94.00 / 11.19 94.33 / 11.01 \$0.39
Ours w/GPT-4 + Golden 98.00 / 8.47 92.67 / 9.87 95.33 / 9.17 \$0.24
Ours w/FT Mistral-7B 77.86 / 11.36 70.47 / 12.04 74.05 / 11.71 -
Table 19: Comparison of turn count and cost across on WebQSP by question type.
Method CWQ
Conj Compo Compa Super Overall Ave Price
Human Anno 1.38 / 12.64 1.10 / 11.32 1.28 / 11.24 1.62 / 11.16 1.34 / 11.59 \$0.28
Ours w/GPT-4 80.67 / 15.25 90.67 / 13.16 72.00 / 15.11 78.67 / 14.75 80.50 / 14.57 \$0.49
Ours w/GPT-4 + Golden 78.00 / 13.55 91.33 / 11.40 75.33 / 13.59 85.33 / 12.44 82.50 / 12.74 \$0.38
Ours w/FT Mistral-7B 49.33 / 14.71 60.00 / 13.69 68.67 / 12.71 80.67 / 12.05 64.67 / 13.29 -
Table 20: Comparison of turn count and cost across on CWQ by question type.
Method KQA Pro
Ct QA QAQ QN QR QRQ
Human Anno 0.86 / 11.28 0.74 / 10.76 1.16 / 14.04 0.96 / 11.12 0.30 / 10.44 0.80 / 14.04
Ours w/GPT-4 96.00 / 11.15 94.00 / 11.57 89.00 / 13.11 86.00 / 10.30 100.00 / 11.00 90.00 / 13.80
Ours w/FT Mistral-7B 94.00 / 10.36 73.00 / 10.70 75.00 / 12.63 64.00 / 10.03 85.00 / 10.20 69.00 / 13.38
(continued) SA SB Vf Overall Ave Price
Human Anno 0.28 / 10.72 0.90 / 12.64 1.20 / 10.68 0.80 / 11.75 \$0.23
Ours w/GPT-4 93.00 / 11.82 99.00 / 13.83 98.00 / 10.61 93.89 / 11.91 \$0.33
Ours w/FT Mistral-7B 83.00 / 10.78 96.97 / 12.35 91.00 / 10.36 81.20 / 11.18 -
Table 21: Comparison of turn count and cost across on KQA Pro by question type.
Method MetaQA
1-hop 2-hop 3-hop Overall Ave Price
Human Anno 0.02 / 6.88 0.12 / 7.22 0.10 / 7.28 0.08 / 7.13 \$0.06
Ours w/GPT-4 99.67 / 7.15 100.00 / 7.14 99.33 / 8.08 99.67 / 7.45 \$0.14
Ours w/FT Mistral-7B 99.00 / 7.01 99.33 / 7.11 98.00 / 7.10 98.78 / 7.07 -
Table 22: Comparison of turn count and cost across on MetaQA by question type.
A.6 Impact of Interaction History Length
To investigate the impact of interaction history length on end-to-end performance, we analyzed the changes in F1 score as the number of dialogue turns increased, as presented in Figures 6. It was observed that as the number of dialogue turns increased, the model’s performance generally deteriorated.
Figure 6: F1 Scores across Different Turn Numbers.
A.7 Confusion Matrix of Classifiers
Figure 7: Confusion matrix of the classifier for CWQ (left) and KQA Pro (right).
Table 7 illustrate the confusion matrices for question type classifiers based on bert-base-uncased.
A.8 Case Study
In Figure 8, 9, 10, and 11, we present four cases. Each illustrates a error and delineates the correction process with human assistance.
Figure 8: Example of a reasoning error: Human assistance in understanding SPARQL logic.
Figure 9: Example of a predicate search error: Human assistance in identifying the predicate.
Figure 10: Example of a hallucination error: Human intervention to rectify the hallucination.
Figure 11: Example of a reasoning error: Human guidance in locating the qualifier.